Find the equation of the quadratic function with zeros 10 and 14 and a vertex at (12, -8)
Answer by scott8148(6628)

the general vertex form is ___ y = a(x - h)^2 + k ___ where (h,k) is the vertex

0 = a(10 - 12)^2 - 8

8 = 4a


y = 2(x - 12)^2 - 8